OpenStudy (anonymous):

antiderivative of -48/x^9

OpenStudy (anonymous):

\[=-48\int\limits_{}^{}x ^{-9}\]

OpenStudy (anonymous):

I got -6x(x^-8) but it's wrong

OpenStudy (anonymous):

\[-48\int\limits_{}^{}\ x ^{-9}=\frac{ x ^{-9+1}}{-9+1}=-48(\frac{ x ^{-8} }{ -8 })=8x ^{-8}\]

OpenStudy (anonymous):

-48 should be in the second term.